OpenStudy (anonymous):

how do you find the product (-1/3x^2-1/3)(1/2x+1/4)

OpenStudy (anonymous):

Do you know how to factor?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

\[\bigg(\frac{1}{3}x^{2}-\frac{1}{3} \bigg) \bigg(\frac{1}{2}x+\frac{1}{4} \bigg)\] Is it right?

OpenStudy (anonymous):

(-1/3x^2-1/3)(1/2x+1/4)

OpenStudy (anonymous):

OK, just the minus in front of first bit, yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\bigg(-\frac{1}{3}x^{2}-\frac{1}{3} \bigg) \bigg(\frac{1}{2}x+\frac{1}{4} \bigg)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

By factor, I mean to take out a common piece from inside a bracket...

OpenStudy (anonymous):

So we could take out the -1/3 from the first bracket for instance.

OpenStudy (anonymous):

and combine it with the 1/2 in the second bracket

OpenStudy (anonymous):

I combined using FOIL and ended up with 1/6x^2 - 1/12 + 1/12

OpenStudy (anonymous):

\[-\frac{1}{3}\bigg (x^{2}+1 \bigg) \bigg(\frac{1}{2}x+\frac{1}{4} \bigg)\] Like this...

OpenStudy (anonymous):

1/6x^2-1/12 +1/6 - 1/12

OpenStudy (anonymous):

I don't understand where u got the above?

OpenStudy (anonymous):

not sure either

OpenStudy (anonymous):

Do you think you could do the same thing as I did with the -1/3, only do it with the second bracket and take out 1/4?

OpenStudy (anonymous):

yes i think I can... I will try and see what happens. thanks for the help

OpenStudy (anonymous):

\[-\frac{1}{3}\bigg (x^{2}+1 \bigg) \dfrac{1}{4} \bigg(2x+1 \bigg)\]

OpenStudy (anonymous):

Now it's easier , right?

OpenStudy (anonymous):

yes it is ... thanks

OpenStudy (anonymous):

ur welcome.